Solving for the pH in a mixture of acids is like dealing with a diprotic acid. You solve the problem by dealing with the acids in successive order. You start with the stronger acid (just like you start with the Ka1 because it is higher than the Ka2 for a diprotic). Afterward, you move on to the weaker acid.

If a solution contains 0.038 M HCl and 0.100 M HF (Ka=6.6x10-4), what will be the pH of the solution?
Answer #1

HCl is a strong acid. Hence, 0.038 M of HCl gives 0.038 M of H3O+ in the solution.

HF is a weak acid. Let us say that it gives X M of H3O+ in the solution on dissociation.

                 HF    +    H2O H3O+ + F-

Initial concentration (M)                                       0.100                         0             0

Change in concentration (M)                                 -X                          X             X

Equilibrium concentration (M)                         (0.100 - X)                     X             X

Given, Ka of HF = 6.6 x 10-4

Now,

Ka = [H3O+][F-]/[HF]

or, 6.6 x 10-4 = (X)(X)/(0.100 -X)

or, 6.6 x 10-4 = (X)(X)/0.100      [(0.100 -X) 0.100 as X << 0.100]

or, X2 = 6.6 x 10-5

or, X = 0.0081

or, [H3O+] = 0.0081 M

Hence, total [H3O+] in the solution = 0.038 M + 0.0081 M

                                                       = 0.0461 M

Now,

pH = - log[H3O+]

      = - log(0.0461)

      = 1.3

Hence, the pH of the solution will be = 1.3